OpenStudy (anonymous):

Find the integral off... (intergrand) 5x cos x^2 dx

OpenStudy (anonymous):

\[\int\limits_{}^{}5xcos^2(x)dx\]

OpenStudy (anonymous):

this is the question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I worked it out but am not sure if it is correct, I dont want to go wolfram it....

OpenStudy (raden):

use u-subst it will work let u=x^2

OpenStudy (anonymous):

wait this is the question \[\int\limits_{0}^{0} x^5 \cos x^2 dx\]

OpenStudy (raden):

oh, which one that question ?

OpenStudy (anonymous):

disregard the 0 on the integrals

OpenStudy (anonymous):

the one I just posted above is the correct one

OpenStudy (anonymous):

I just fully mucked that up. I apologize everyone. This is the question \[\int\limits_{?}^{?}5x \cos x^2 dx\]

OpenStudy (anonymous):

Im a dumbass so sorry about this

OpenStudy (raden):

but, for integral f(x) with interval from a to a, it can be zero... right, @Coolsector ?

OpenStudy (anonymous):

disregard the zeros on the integrals, I didnt know how to remove them

OpenStudy (anonymous):

i think he wants indefinite integral

OpenStudy (anonymous):

use u sub as RadEn suggested

OpenStudy (raden):

ohhhhhh, lol

OpenStudy (anonymous):

u=x^2 then du=2xdx

OpenStudy (anonymous):

then your integral becomes 2.5 cos(u)du

OpenStudy (anonymous):

the cos will turn into sin?

OpenStudy (anonymous):

yes integrating this will give you 2.5sin(u) which is 2.5sin(x^2)

OpenStudy (anonymous):

you can verify your answer by differentiating 2.5sin(x^2) and getting 5xcos(x^2) so it is correct :)

OpenStudy (anonymous):

I got without substitution \[f(x) =\frac{ 5 }{ 2 } \sin x^2\]

OpenStudy (anonymous):

it is good ..

OpenStudy (anonymous):

ok great, thank you guys

OpenStudy (anonymous):

dont forget constant of integration +C ;)

OpenStudy (anonymous):

yes thats right, my teachers nemonic for that is cough, so you need vit c to boost your imune lol

OpenStudy (calculusfunctions):

Yes but it still deserves a 0 on a test because you forgot to add a constant c.

OpenStudy (raden):

so, without intervals ?

OpenStudy (anonymous):

LOL

OpenStudy (calculusfunctions):

Without intervals it's called an indefinite integral.

OpenStudy (anonymous):

ok thank you guys.